OpenStudy (anonymous):

Tammy works as a tutor for $7 an hour and as a waitress for $11 an hour. This month, she worked a combined total of 98 hours at her two jobs. Let t be the number of hours Tammy worked as a tutor this month. Write an expression for the combined total dollar amount she earned this month.

ganeshie8 (ganeshie8):

names and numbers have changed essentially this is same as the previous problem

OpenStudy (anonymous):

7t+11(98-t)

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

Let n be the middle number of three consecutive integers. Write an expression for the sum of these integers.

ganeshie8 (ganeshie8):

start by writing expressions for immediate left and right integers : \[\large \_ , ~n, ~\_\]

OpenStudy (anonymous):

2,n,8

ganeshie8 (ganeshie8):

interesting, how did u get 2 and 8 ?

ganeshie8 (ganeshie8):

you need to write them in terms of \(n\)

OpenStudy (anonymous):

just guessing

OpenStudy (anonymous):

how

ganeshie8 (ganeshie8):

consecutive integers are the integers that follow each other in order for example : 3,4,5 are consecutive

OpenStudy (anonymous):

(n-1)+n+(n+1)

ganeshie8 (ganeshie8):

Brilliant !

OpenStudy (anonymous):

really?

ganeshie8 (ganeshie8):

yes thats it!

ganeshie8 (ganeshie8):

if you're not sure just plugin some value for \(n\) and verify if you really get consecutive integers

OpenStudy (anonymous):

Translate the sentence into an equation. Four times the sum of a number and 8 equals 2 . Use the variable c for the unknown number.

ganeshie8 (ganeshie8):

should be easy for you, give it a try

OpenStudy (anonymous):

4c-8=2

ganeshie8 (ganeshie8):

Four times the `sum` of a number and 8 equals 2

ganeshie8 (ganeshie8):

forget about Four times for the time being how do you write the expression for "sum of c and 8" ?

OpenStudy (anonymous):

SORRY ITS 2:33 over here

OpenStudy (anonymous):

am

ganeshie8 (ganeshie8):

its okay..

ganeshie8 (ganeshie8):

you should be sleeping lol

ganeshie8 (ganeshie8):

Four times the sum of a number and 8 equals 2 break the senternce like this : `Four times` the `sum of a number and 8` equals `2`

ganeshie8 (ganeshie8):

\[4(c+8) = 2\]

OpenStudy (anonymous):

ok we find c or thats it

ganeshie8 (ganeshie8):

question asks u only to translate the given sentence so we're good

OpenStudy (anonymous):

Translate the sentence into an equation. Two less than the product of 6 and a number is 3 . Use the variable c for the unknown number.

OpenStudy (anonymous):

let me try

ganeshie8 (ganeshie8):

kk

OpenStudy (anonymous):

2-c+2=3

ganeshie8 (ganeshie8):

it is slightly tricky compared to previous one

OpenStudy (anonymous):

well i tryed lol

OpenStudy (anonymous):

-2(c-2)=3

ganeshie8 (ganeshie8):

thats even better, lets breal the given sentence : `Two less than` the `product of 6 and a number` is `3` .

ganeshie8 (ganeshie8):

`-2` `6c` `3`

ganeshie8 (ganeshie8):

\[6c - 2 =3\]

OpenStudy (anonymous):

kk i have more bt putting it on another page

ganeshie8 (ganeshie8):

Okay sure one question per thread would be good as the new questions get bumped up automatically